OpenStudy (anonymous):

Please, please help with substitution rule for indefinite integrals:

OpenStudy (anonymous):

\[\int\limits_{}^{} (4z + \sin(3z))e ^{6z ^{2}-\cos(3z)} dz\]

OpenStudy (anonymous):

whew, that's a doozy.

OpenStudy (anonymous):

So what do you think you want to use for substitution?

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \sin(x) d x = -\cos(x) + constant\]

OpenStudy (anonymous):

Not sure...possibly something to do with 3z? I have no idea to be honest

OpenStudy (anonymous):

I would let u = that yucky exponent on the e

OpenStudy (anonymous):

Ignore the two ?'s

OpenStudy (anonymous):

Okay but how does that help?

OpenStudy (anonymous):

Well. \[u = 6x^2 - cos(3z)\] \[\implies du = 12z +3sin(3z)dz = 3(4z + sin(3z))dz\]

OpenStudy (anonymous):

Oh so then \[du = 2z ^{3}-\sin(3z)dx\]

OpenStudy (anonymous):

Close, but you integrated. You want to take the derivative

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Okay. Where do I plug u into?

OpenStudy (anonymous):

Well first rewrite dz in terms of du using that last equation I had.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\(du = 3(4z + sin(3z))dz \implies dz =\ ?\)

OpenStudy (anonymous):

I don't understand - what am I solving for?

OpenStudy (anonymous):

Well for this part you're solving for dz.

OpenStudy (anonymous):

Err...I divide 3(4z + sin(3z)) on both sides?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So \[dz = \frac{1}{3(4z+sin(3z))} du\]

OpenStudy (anonymous):

So now. Go back to your integral, and replace the exponent of e with u, and the dz with your expression for du, and you'll be pleasantly surprised.

OpenStudy (anonymous):

Wait a minute...Why do I replace dz?

OpenStudy (anonymous):

Because we are no longer integrating with respect to z, we are integrating with respect to u now. Because it turns out u is nicer.

OpenStudy (anonymous):

We are replacing our nasty expression with z for a nice expression with u.

OpenStudy (anonymous):

\[(4z + \sin(3z)e ^{u}\]du

OpenStudy (anonymous):

But remember that dz does not equal du.

OpenStudy (anonymous):

ButI thought we replaced dz

OpenStudy (anonymous):

\[dz = \frac{1}{3(4z+sin(3z))} du\]

OpenStudy (anonymous):

So you have to put that whole thing in for dz

OpenStudy (anonymous):

I'm so lost

OpenStudy (anonymous):

Lets try an easy one to start with.

OpenStudy (anonymous):

\[\int x*e^{x^2}dx\]

OpenStudy (anonymous):

We let \(u = x^2\)

OpenStudy (anonymous):

So \[du = (2x)dx \implies dx = \frac{1}{2x}du\]

OpenStudy (anonymous):

With me so far?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

If not, please ask

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok. So then \[\int x*e^{x^2}dx = \int x*e^u(\frac{1}{2x} du)\]

OpenStudy (anonymous):

And we can cancel the x in front with the x in the denominator from the du.

OpenStudy (anonymous):

to get \[\frac{1}{2}\int e^udu\]

OpenStudy (anonymous):

Which is nice and easy.

OpenStudy (anonymous):

\[\frac{1}{2}\int e^udu = \frac{1}{2}e^u + c = \frac{1}{2}e^{x^2} + C\]

OpenStudy (anonymous):

So I've got\[1/3 e ^{u}\] for my problem when do I integrate?

OpenStudy (anonymous):

Now.

OpenStudy (anonymous):

Before or after I plug in u

OpenStudy (anonymous):

Don't switch back your u until after you integrate.

OpenStudy (anonymous):

How do I integrate that?

OpenStudy (anonymous):

\[\int \frac{1}{3}e^u du = \frac{1}{3}\int e^udu\]

OpenStudy (anonymous):

What is the integral of \(e^x\)?

OpenStudy (anonymous):

e^x

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

\[1/3 *e ^{6z ^{2}-\cos(3z)} \]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

Nice job!

OpenStudy (anonymous):

Oh my goodness. Would you please walk me through one more?

OpenStudy (anonymous):

Sure, but you'll have to do more of it this time ;)

OpenStudy (anonymous):

How's it start?

OpenStudy (anonymous):

Okay I'll try:

OpenStudy (anonymous):

\[\int\limits_{}^{}(4z ^{3} + 6\csc ^{2}(6z)(z ^{4}-\cot(6z))^{4} dz\]

OpenStudy (anonymous):

Ugh

OpenStudy (anonymous):

Do I need to FOIL first?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

These are all ugly u substitution problems. Foiling will just make you cry.

OpenStudy (anonymous):

You want to find a nice thing to pick for a u substitution.

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Do you notice anything interesting about these two things?

OpenStudy (anonymous):

Not really sure what to look for, actually.

OpenStudy (anonymous):

\((4z^3 + 6csc^2(6z)) * (z^4 - cot(6z))^4\) You are looking for something in the equation that looks like the derivative of another part of the equation.

OpenStudy (anonymous):

Err expression, not equation

OpenStudy (anonymous):

Oh! I guess the 4z^3 and z^4

OpenStudy (anonymous):

And what about the \(6csc^2(6z)\) and the -cot(6z) ?

OpenStudy (anonymous):

AND the csc^2

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

So what you want to do is pick the one that is not the derivative, and let that be u.

OpenStudy (anonymous):

So let\[u = z ^{4} - \cot(6z)\]

OpenStudy (anonymous):

Yep, then solve for dz in terms of z and du.

OpenStudy (anonymous):

So \[\int\limits_{}^{}(4z ^{3}+6\csc ^{2}(6z))u ^{4}du\]

OpenStudy (anonymous):

Nope. but close.

OpenStudy (anonymous):

\[u = z^4 - cot(6z)\] So \[du =\ ?\]

OpenStudy (anonymous):

4z^3 + 6csc^2(6z)

OpenStudy (anonymous):

close.

OpenStudy (anonymous):

\[du = (4z^3 + 6csc^2(6z))dz\]

OpenStudy (anonymous):

So then what does dz = ?

OpenStudy (anonymous):

1 over all that mess

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

So plug in 1 over all that mess times du for dz in the original equation. and plug in u for what we said it equaled.

OpenStudy (anonymous):

\[(4z ^{3} + 6\csc ^{2}(6z))u ^{4} * 1/4z ^{3} + 6\csc ^{2}(6z)\]

OpenStudy (anonymous):

simplify!

OpenStudy (anonymous):

u^4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(z ^{4} - \cot(6z))^{4}\]

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[\int(4z^3 + 6csc^2(6z)) * (z^4 - cot(6z))^4dz\] \[ = \int (4z^3 + 6csc^2(6z)) u^4(\frac{1}{4z^3 + 6csc^2(6z)})du \] \[= \int u^4du =\ ?\]

OpenStudy (anonymous):

? I don't know

OpenStudy (anonymous):

I can't pull anything out in front of the integral sign